### v1

@@ A.1 Fourier Analysis over $\mathbb{Z}_{p_{1}}^{n_{1}}\times\dots\times\mathbb{Z}_{p_{t}}^{n_{t}}$
Proof A.13.
$|\widehat{fg}(\chi)|$ $=|\sum_{\psi\in\widehat{\mathbb{Z}_{p}^{n}}}\widehat{f}(\psi)\widehat{g}(\chi+\psi)|$
[…]
@@
Definition A.14 ($\mu$-close to $s$-sparse).
Let $f$ and $g$ be two functions with domain $\mathbb{Z}_{p}^{n}$ and range $\mathbb{C}$. Then the square of the $\ell_{2}$ distance between $f$ and $g$ is defined as $\mathbb{E}_{x\in\mathcal{G}}[|f(x)-g(x)|^{2}].$ By Parseval’s identity the square of the $\ell_{2}$-distance between $f$ and $g$ can also be written as $\sum_{\chi\in\widehat{\mathbb{Z}_{p}^{n}}}|\widehat{(f-g)}(\chi)|^{2}$.
@@
The following lemma gives us a relation between the $\ell_{2}$ distance between two Boolean valued functions $f$ and $g$ defined in Definition A.14 and $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$.
Lemma A.15.
The square of the $\ell_{2}$ distance between two Boolean valued functions $f$ and $g$ defined in Definition A.14 is equal to $4\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$.  [^2] If $f$ and $g$ are two Booelan-valued functions then $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$ is also called the Hamming distance between the two functions. So the $\ell_{2}$ norm between two Boolean-valued functions is 4 times the Hamming distance between two Boolean-valued functions.
Proof A.16.
By Parseval (Theorem A.8), we have,
@@
So, the square of the $\ell_{2}$ distance between $f$ and $g$ is equal to $4$ times $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$.
@@
$||P-Q||_{TV}=\frac{1}{2}\sum_{x}|P(x)-Q(x)|.$
A.2 Basics from Probability
Definition A.19 (Covariance).
Lemma A.20 (Markov’s inequality).
Let $X$ be a nonnegative random variable and $a>0$, then,
$\Pr[X\geq a]\leq\frac{\mathbb{E}[X]}{a},$
where $\mathbb{E}[X]$ is the expected value of $X$.
Lemma A.21.
  1. 1.
    $\Pr[X\leq(1-\delta)\mathbb{E}[X]]\leq e^{-\frac{\delta^{2}\mathbb{E}[X]}{2}}$, $10<\delta<$ 1.
  2. 2.
Lemma A.22 (Hoeffding’s Inequality).
A.3 Basics from algebraic number theory
Definition A.23.
Recall that $\alpha\in\mathbb{C}$ is an algebraic number if it is the root of a non-zero polynomial in $\mathbb{Q}[x]$. The minimal polynomial of $\alpha$ is the unique monic polynomial in $\mathbb{Q}[x]$ having $\alpha$ as a root. If the minimal polynomial has integer coefficients, then we say that $\alpha$ is an algebraic integer. It is a well-known fact that a cyclotomic integer is an algebraic integer. For details, see [23, Chapter 13], or expository lecture notes [16, Chapter 2] & [30].
Recall that the Galois group, denoted $\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})$ of automorphisms of $Q(\omega_{n})$; it is isomorphic to the multiplicative group $\mathbb{Z}_{n}^{\ast}$, of integers $\bmod\leavevmode\nobreak\ n$, i.e., all integers $k$ such that gcd$(k,n)=1$. Note that $|\mathbb{Z}_{n}^{\ast}|=\varphi(n)$, where $\varphi$ is the Euler’s totient function, where $\varphi(n):=n\cdot\prod_{\text{prime}\leavevmode\nobreak\ p\mid n}(1-1/p)$. In simple words, think of $\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})$, as consisting $\varphi(n)$ many automorphism $\sigma_{k}$, that sends $\omega_{n}\mapsto\omega_{n}^{k}$, where $k\in\mathbb{Z}_{n}^{\ast}$. For basics, we refer to [16, Chapter 1].
Definition A.24 (Norm).
It is a well-known fact that for a non-zero $\alpha$, $N_{\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q}}(\alpha)\in\mathbb{Q}-\{0\}$; see [16, Lemma 1.6.1]. In particular, when $\alpha$ is a cyclotomic integer, i.e, $\alpha\in\mathbb{Z}[\omega_{n}]$, it is well-known that $N_{\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q}}(\alpha)\in\mathbb{Z}$, see [16, Corollary 2.2.2]. We restate the same as a theorem below in the most simplified way without the above terminologies since this fact has been crucially used throughout the paper.
Theorem A.25.
@@ A.3 Basics from algebraic number theory
In retrospect, one can think of $\alpha=f(\omega_{n})\in\mathbb{Z}[\omega_{n}]$. Note that, $\sigma_{k}(\alpha)=f(\omega_{n}^{k})$, by the definition of the automorphism $\sigma:\omega_{n}\mapsto\omega_{n}^{k}$, where gcd$(k,n)=1$. Therefore, their product, which is the norm $N_{\mathsf{Gal}(\mathbb{Q}(\omega_{n})/\mathbb{Q})}(\alpha)$, must be a non-zero integer, and hence the conclusion follows. Using Theorem 2.1, one can directly conclude the following fact (which can also be proved in an elementary way).
Fact 1.
Let $f\in\mathbb{Z}[x]$, such that $f(\omega_{p})\neq 0$. Then, $f(\omega_{p}^{i})\neq 0$, for any $i\in[p-1]$.
Appendix B Proofs of structure Theorems 1.3 and  1.4
In this section, we will prove Theorem 1.3 and Theorem 1.4 which are generalizations of Gopalan et al’s result  [15] to the case when the domain of the function is $\mathbb{Z}_{p}^{n}$, where $p$ is an odd prime. Let $\omega_{p}$ be a primitive $p^{th}$ root of unity. For any polynomial $\sum_{i=0}^{d}g_{i}x^{i}=g(x)\in\mathbb{C}[x]$, let $|g|_{1}\;:=\;\sum_{i=0}^{d}|g_{i}|$.
The proof of the more general result of Theorem 1.5 has been relegated to the Appendix F.
B.1 Proof of Theorem 1.3
In this subsection we will prove Theorem 1.3. Recall the definition of granularity for complex numbers (Definition 1.1) and the definition of $\mu$-close to granular (Definition 1.2). Before moving forward, we need the following definition.
Definition B.1.
(Projection operator $P_{r+H}$) Given a subgroup $H=H_{1}\times\cdots\times H_{t}\subseteq\mathcal{G}$ and a coset $r+H$, where for each $i$, $H_{i}$ is a subgroup of $\mathbb{Z}_{p_{i}}^{n_{i}}$ (hence a subspace), let us define the projection operator $P_{r+H}$ on functions $f:\mathcal{G}\to\mathbb{R}$ as follows:
@@
Lemma B.2.
Given a subspace $H$ of $\mathbb{Z}_{p}^{n}$, we have
@@
Proof B.3 (Proof of Lemma B.2).
From the definition of the projection operator $P_{r+H}$ (Definition B.1) we get
$P_{r+H}f(x)$ $=\sum_{\alpha\in r+H}\hat{f}(\alpha)\chi_{\alpha}(x)$
[…]
@@
$=\chi_{r}(x)\sum_{\beta\in H}\chi_{\beta}(x)\frac{1}{|\mathbb{Z}_{p}^{n}|}\sum_{y\in\mathbb{Z}_{p}^{n}}f(y)\overline{\chi_{r+\beta}(y)}$
$=\frac{1}{|\mathbb{Z}_{p}^{n}|}\chi_{r}(x)\sum_{y\in\mathbb{Z}_{p}^{n}}f(y)\sum_{\beta\in H}\chi_{\beta}(x)\chi_{-r-\beta}(y)$
$=\frac{1}{|\mathbb{Z}_{p}^{n}|}\chi_{r}(x)\sum_{y\in\mathbb{Z}_{p}^{n}}f(y)\sum_{\beta\in H}\chi_{\beta}(x)\chi_{r+\beta}(-y)$
$=\frac{1}{|\mathbb{Z}_{p}^{n}|}\chi_{r}(x)\sum_{y\in\mathbb{Z}_{p}^{n}}f(y)\chi_{r}(-y)\sum_{\beta\in H}\chi_{\beta}(x-y)$
$=\frac{1}{|\mathbb{Z}_{p}^{n}|}\chi_{r}(x)\sum_{y\in\mathbb{Z}_{p}^{n}}f(y)\chi_{r}(-y)|H|1_{H^{\perp}}(x-y)$
$=\frac{1}{|H^{\perp}|}\chi_{r}(x)\sum_{x-y\in H^{\perp}}f(y)\chi_{r}(-y)$
$=\frac{1}{|H^{\perp}|}\sum_{x-y\in H^{\perp}}f(y)\chi_{r}(x-y)$
[…]
$=\mathbb{E}_{z\in H^{\perp}}[f(x-z)\chi_{r}(z)].$
Now we prove Theorem 1.3.
Proof B.4 (Proof of Theorem 1.3).
Set $k=\lceil\log_{p}s\rceil+1$, and let $A$ be a $k\times n$ matrix and $b$ be a $k$-dimensional vector.
Let $A^{\perp}\subset\mathbb{Z}^{n}_{p}$, be the set of solutions to $A\chi=0$. Let $H$ be the coset of $A^{\perp}$ that are solutions to system of linear equations $A\chi=b$. From the definition,
$P_{H}f(x)\;=\;\sum_{\chi\in H}\widehat{f}(\chi)\chi(x).$ (B.1)
We need to show that for each $\chi_{r_{i}}\in B$, there exists a $k\times n$ matrix $A$ and a $k$-dimensional vector $b$ such that
$\{\chi\,|\,\chi\in B\}\bigcap\left\{\chi\,|\,A\chi\,=\,b\right\}\;=\;\{\chi_{r_{i}}\},$ (B.2)
@@
To show the existence of $A$ and $b$, given $\chi_{r_{i}}\in B$, for which Equation B.2 is true we will need to consider two cases:
@@
 [^3] Here $\chi_{r_{i}}=0$ means that $\chi_{r_{i}}=\chi_{r}$ where $r=(0,\dots,0)\in\mathbb{Z}^{n}_{p}$.
. Observe that for all $A$, we have $A\chi_{r_{i}}=0$. If entries of $A$ are sampled independently and uniformly at random from $\mathbb{Z}_{p}$. Then, for all $\chi_{r_{j}}\in B\setminus\{\chi_{r_{i}}\}$, we have
$\Pr_{A}\left[A\chi_{r_{j}}=0\right]=\frac{1}{p^{k}}\leq\frac{1}{s}.$
Note that $\Pr_{A}\left[E_{1}\right]$ denotes the probability of event $E_{1}$ when $A$ is the random matrix generated by the above-mentioned procedure. Using union bound, we get
$\Pr_{A}\left[\exists j\neq i\text{ such that }A\chi_{r_{j}}=0\right]\leq\frac{s-1}{s}\leq\frac{p^{k-1}-1}{p^{k}}<1$
Therefore, if $\chi_{r_{i}}=0$ then there exists a $k\times n$ matrix $A$ such that
[…]
@@
 [^4] Here $\chi_{r_{i}}=0$ means that $\chi_{r_{i}}=\chi_{r}$ where $r\neq(0,\dots,0)\in\mathbb{Z}^{n}_{p}$.
. Fix a nonzero vector $b\in\mathbb{Z}^{n}_{p}\setminus\{(0,\dots,0)\}$. Again, let $A$ be a random $k\times n$ matrix with entries of $A$ being sampled independently and uniformly at random from $\mathbb{Z}_{p}$. Observe that $\Pr_{A}[A\chi_{r_{i}}=b]=\frac{1}{p^{k}}$. Let $\chi_{r_{j}}\in B\setminus\{\chi_{r_{i}}\}$. If $\chi_{r_{j}}=\alpha\chi_{r_{i}}$, where $\alpha\in\mathbb{Z}_{p}\setminus\{1\}$, then
$\Pr_{A}[A\chi_{r_{j}}\,=\,b\,\mid\,A\chi_{r_{i}}\,=\,b]=0.$
Now we will consider the case when $\chi_{r_{j}}$ is not a scalar multiple of $\chi_{r_{i}}$. Observe that
$\Pr_{A}[A\chi_{r_{j}}\,=\,b\,\mid\,A\chi_{r_{i}}\,=\,b]$ $=\frac{\Pr_{A,b}[A\chi_{r_{j}}\,=\,b,A\chi_{r_{i}}\,=\,b]}{\Pr_{A}[A\chi_{r_{i}}\,=\,b]}$
$=\frac{\Pr_{A}\left[\bigcap\limits_{\ell\in[k]}\left(A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell}\right)\right]}{\Pr_{A}[A\chi_{r_{i}}\,=\,b]}$
$=\frac{\prod\limits_{\ell\in[k]}\Pr_{A}\left[A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell}\right]}{\Pr_{A}[A\chi_{r_{i}}\,=\,b]}$ (B.3)
Note that $A_{\ell}$ and $b_{\ell}$ denote the $\ell$-th row of $A$ and $\ell$-th coordinate of $b$ respectively, and $[k]$ denotes the set $\{1,\dots,k\}$. Observe that the last equality in the above equation follows from the fact that rows of $A$ are independent and uniformly random samples of $\mathbb{Z}^{n}_{p}$. As $\chi_{r_{j}}$ and $\chi_{r_{i}}$ are not scalar multiple of each other therefore
$\Pr_{A}\left[A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell}\right]=\frac{1}{p^{2}}.$
Using the above equation and the fact that $\Pr_{A}[A\chi_{r_{i}}=b]=\frac{1}{p^{k}}$ in Equation (Case 2: $\chi_{r_{i}}\neq 0$) we get
$\Pr_{A}[A\chi_{r_{j}}\,=\,b\,\mid\,A\chi_{r_{i}}\,=\,b]=\frac{\prod\limits_{\ell\in[k]}\Pr_{A}\left[A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell}\right]}{\Pr_{A}[A\chi_{r_{i}}\,=\,b]}=\frac{\left(\frac{1}{p^{2}}\right)^{k}}{\frac{1}{p^{k}}}=\frac{1}{p^{k}}$
@@
$\Pr_{A}[\exists j\neq i\text{ such that }A\chi_{r_{j}}=b\mid A\chi_{r_{i}}=b]\leq\frac{s-1}{p^{k}}\leq\frac{p^{k-1}-1}{p^{k}}<1.$
From the above two cases we can now say that given any $\chi_{r_{i}}\in B$, there exists a $k\times n$ matrix $A$ and a $k$-dimensional vector $b$ such that
@@
Now, let us look at the coefficients corresponding to the characters in $S$, in order to prove that the $\ell$ Fourier weight of $S\cap H$ is upper bounded by $\frac{\mu^{2}}{s}$, $H$ being the solution set of the system of linear equations $A\chi=b$. Let us look at the expected value of $\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}$, given $A\chi_{r_{i}}=b$.
$\mathbb{E}_{A,b}\biggl{[}\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}\mid A\chi_{r_{i}}=b\biggr{]}=\sum_{\chi\in S}\Pr_{A,b}[\chi\in H\mid A\chi_{r_{i}}=b]|\widehat{f}(\chi)|^{2}\leq\frac{\mu^{2}}{p^{k}}\leq\frac{\mu^{2}}{ps},$ (B.4)
since $k=\lceil\log_{p}s\rceil+1$.
@@
$\Pr_{A,b}\bigg{[}\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}\geq\frac{\mu^{2}}{s}\mid A\chi_{r_{i}}=b\bigg{]}\leq\frac{\frac{\mu^{2}}{ps}}{\frac{\mu^{2}}{s}}=\frac{1}{p}<1.$ (B.5)
So, if we apply union bound to Equation B.2 and Equation B.5, we get that the events $E_{1}$ and $E_{2}$ occur simultaneously with non-zero probability.
@@
Now, by Lemma B.2, we have,
$P_{H}f(x)=\mathbb{E}_{z\in r(A)}[f(x-z)\chi_{r}(z)],$
which is of the form $\frac{g(\omega_{p})}{p^{k}}$, where $g(X)\in\mathbb{Z}[X]$ is a nonzero polynomial with degree $\leq p-1$, and $r(A)$ is the row-space of $A$. So $P_{H}f(x)$ is $k$-granular.
Also, let us look at the function $G_{1}(x)=\sum_{\chi\in S\cap H}\widehat{f}(\chi)\chi(x)$. We know, by Equation B.4, that, $\mathbb{E}_{x}G_{1}(x)^{2}\leq\frac{\mu^{2}}{s}$. So, $G_{1}(x)^{2}\leq\frac{\mu^{2}}{s}$ for some $x\in\mathbb{Z}_{p}^{n}$. For this particular $x$,

### v2

@@ 2.5 Miscellaneous
Proof.
$|\widehat{fg}(\chi)|$ $=|\sum_{\psi\in\widehat{\mathcal{G}}}\widehat{f}(\psi)\widehat{g}(\chi+\psi)|$
$\leq\sqrt{\sum_{\psi\in\widehat{\mathcal{G}}}\widehat{f}(\psi)^{2}}\sqrt{\sum_{\psi\in\widehat{\mathcal{G}}}\widehat{g}(\chi+\psi)^{2}}\text{ by Cauchy-Schwartz inequality}$
@@
Definition 2.42 ($\mu$-close to $s$-sparse).
Let $f$ and $g$ be two functions with domain $\mathcal{G}$ and range $\mathbb{C}$. Then the square of the $\ell_{2}$ distance between $f$ and $g$ is defined as $\mathbb{E}_{x\in\mathcal{G}}[|f(x)-g(x)|^{2}].$ By Parseval’s identity the square of the $\ell_{2}$-distance between $f$ and $g$ can also be written as $\sum_{\chi\in\widehat{\mathcal{G}}}|\widehat{(f-g)}(\chi)|^{2}$.
@@
The following lemma gives us a relation between the $\ell_{2}$ distance between two Boolean valued functions $f$ and $g$ defined in Definition 2.42 and $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$.
Lemma 2.43.
The square of the $\ell_{2}$ distance between two Boolean valued functions $f$ and $g$ defined in Definition 2.42 is equal to $4\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$.  [^6] If $f$ and $g$ are two Booelan-valued functions then $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$ is also called the Hamming distance between the two functions. So the $\ell_{2}$ norm between two Boolean-valued functions is 4 times the Hamming distance between two Boolean-valued functions.
Proof.
By Parseval (Theorem 2.10), we have,
@@
So, the square of the $\ell_{2}$ distance between $f$ and $g$ is equal to 44 times $\Pr_{x}[x\in\mathcal{G}|f(x)\neq g(x)]$. ∎
@@
3 Proofs of structure Theorems 1.3 and  1.4
In this section, we will prove Theorem 1.3 and Theorem 1.4 which are generalizations of Gopalan et al’s result  [GOS+11] to the case when the domain of the function is $\mathcal{G}$. Note that Chakraborty et al [CDD+24] proved these results for groups of the form $\mathbb{Z}_{p_{1}}^{n_{1}}\times\cdots\times\mathbb{Z}_{p_{T}}^{n_{T}}$, where $p_{i},i\in[T]$ are distinct primes. Since $\mathcal{G}$ is not a vector space in general, we use the pseudo-inner product, normal subgroups, and cosets (see Section 2 for more details) to prove our results. Also, let $\omega_{L}$ be a primitive $L^{th}$ root of unity, $L$ is a positive integer. For any polynomial $\sum_{i=0}^{d}g_{i}x^{i}=g(x)\in\mathbb{C}[x]$, let $|g|_{1}\;:=\;\sum_{i=0}^{d}|g_{i}|$.
3.1 The projection on a coset
To prove Theorem 1.3, we need the following. Here $\mathcal{G}$ denotes a finite Abelian group.
Definition 3.1.
(Projection $P_{r+H}$) Given a subgroup $H\subseteq\mathcal{G}$ and a coset $r+H$, let us define the projection operator $P_{r+H}$ on functions $f:\mathcal{G}\to\mathbb{R}$ as follows:
@@ 3.1 The projection on a coset
Lemma 3.2.
Given a subgroup $H$ of $\mathcal{G}$, we have
@@
Proof of Lemma 3.2.
From the definition of the projection operator $P_{r+H}$ (Definition 3.1) we get
Proof.
$\mathcal{P}_{r+H}f(x)$ $=\sum_{\gamma\in r+H}\hat{f}(\gamma)\chi_{\gamma}(x)$
$=\sum_{\beta\in H}\hat{f}(r+\beta)\chi_{r+\beta}(x)$ $\text{ where }\gamma=r+\beta,\beta\in H$
@@
$=\chi_{r}(x)\sum_{\beta\in H}\chi_{\beta}(x)\frac{1}{|\mathcal{G}|}\sum_{y\in\mathcal{G}}f(y)\overline{\chi_{r+\beta}(y)}$
$=\frac{1}{|\mathcal{G}|}\chi_{r}(x)\sum_{y\in\mathcal{G}}f(y)\sum_{\beta\in H}\chi_{\beta}(x)\chi_{-r-\beta}(y)$
$=\frac{1}{|\mathcal{G}|}\chi_{r}(x)\sum_{y\in\mathcal{G}}f(y)\sum_{\beta\in H}\chi_{\beta}(x)\chi_{r+\beta}(-y)$
$=\frac{1}{|\mathcal{G}|}\chi_{r}(x)\sum_{y\in\mathcal{G}}f(y)\chi_{r}(-y)\sum_{\beta\in H}\chi_{\beta}(x-y)$
$=\frac{|H|}{|\mathcal{G}|}\sum_{y\in H^{\perp}}f(y)\chi_{r}(x-y)$ by 2.25
$=\frac{1}{|H^{\perp}|}\sum_{z\in H^{\perp}}f(x-z)\chi_{r}(z)$ by Lemma2.27
$=\mathbb{E}_{z\in H^{\perp}}[f(x-z)\chi_{r}(z)],$
where z=xyz=x-y. ∎
3.2 Proof of Theorem 1.3
In this subsection we will prove Theorem 1.3. All Recall the definition of granularity for complex numbers (Definition 1.1) and the definition of $\mu$-close to granular (Definition 1.2).
Proof of Theorem 1.3.
Set $k=\lceil\log_{\Delta}2s\rceil$, and let AA be a $k\times t$ matrix with the entries $a_{ij}$ such that $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$, and $b$ be a $k\times 1$ matrix.
Let $E$ be the event that the matrix AA is chosen such that its entries $a_{ij}$ are invertible in $\mathbb{Z}_{\Delta}$ for all $i\geq j$, and $a_{ij}$ are chosen independently and uniformly at random.
Let $A^{\perp}\subset\mathcal{G}$, be the set of solutions to $A\chi=0$. Let $H$ be the coset of $A^{\perp}$ that are solutions to system of linear equations $A\chi=b$. From the definition,
$P_{H}f(x)\;=\;\sum_{\chi\in H}\widehat{f}(\chi)\chi(x).$ (3.3)
We need to show that for each $\chi_{r_{i}}\in B$, there exists a $k\times n$ matrix AA, with the entries $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$, and a $k\times 1$ matrix $b$ such that
$\{\chi\,|\,\chi\in B\}\bigcap\left\{\chi\,|\,A\chi\,=\,b\right\}\;=\;\{\chi_{r_{i}}\},$ (3.4)
@@
To show the existence of AA and $b$, given $\chi_{r_{i}}\in B$, for which Eq. 3.4 is true we will need to consider two cases:
@@
 [^7] Here $\chi_{r_{i}}=0$ means that $\chi_{r_{i}}=\chi_{r}$ where $r=(0,\dots,0)\in\mathcal{G}$.
. Observe that for all AA, we have $A\chi_{r_{i}}=0$, so $\Pr_{A}[A\chi_{r_{i}}=0]=1$. If entries of AA are sampled independently and uniformly at random from $\mathbb{Z}_{\Delta}$, with $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$, then, for all $\chi_{r_{j}}\in B\setminus\{\chi_{r_{i}}\}$, we have
$\Pr_{A}\left[A\chi_{r_{j}}=0\mid A\chi_{r_{i}}=0,E\right]=\frac{\Pr_{A}\left[\bigl{(}A\chi_{r_{j}}=0,A\chi_{r_{i}}=0\bigr{)},\ E\right]}{\Pr_{A}\left[\bigl{(}A\chi_{r_{i}}=0\bigr{)},\ E\right]}=\frac{1}{\Delta^{k}}.$
Note that $\Pr_{A}\left[E_{1}\right]$ denotes the probability of event $E_{1}$ when AA is the random matrix generated by the above-mentioned procedure. Using union bound, we get
$\Pr_{A}\left[\biggl{(}\exists j\neq i\text{ such that }A\chi_{r_{j}}=0\biggr{)}\mid E\right]=\frac{s-1}{2s}<1.$
Therefore, if $\chi_{r_{i}}=0$ then there exists a $k\times t$ matrix AA, with the entries $a_{ij}$ are invertible for all $i\geq j$ such that
$\{\chi\,|\,\chi\in B\}\bigcap\left\{\chi\,|\,A\chi\,=\,0\right\}=\{\chi_{r_{i}}\}.$
@@
 [^8] Here $\chi_{r_{i}}\neq 0$ means that $\chi_{r_{i}}=\chi_{r}$ where $r\neq(0,\dots,0)\in\mathcal{G}$.
. Fix a nonzero element $b\in\mathcal{G}\setminus\{(0,\dots,0)\}$. Again, let AA be a random $k\times t$ matrix with entries of AA being sampled independently and uniformly at random from $\mathbb{Z}_{\Delta}$ and the entries $a_{ij}$ are invertible in $\mathbb{Z}_{\Delta}$ for all $i\geq j$. If $\chi_{r_{j}}=\alpha\chi_{r_{i}}$, where $\alpha\in\mathbb{Z}_{\Delta}$ and $\gcd(\alpha-1,\Delta)=1$, then
$\Pr_{A}[A\chi_{r_{j}}\,=\,b\,\mid\,A\chi_{r_{i}}\,=\,b,\ E]=0.$
Now we will consider the case when $\chi_{r_{j}}\neq\alpha\chi_{r_{i}}$, with $\alpha\in\mathbb{Z}_{\Delta}$ and $\gcd(\alpha-1,\Delta)=1$. Observe that
$\Pr_{A}[A\chi_{r_{j}}\,=\,b\,\mid\,A\chi_{r_{i}}\,=\,b,\ E]$ $=\frac{\Pr_{A,b}[\bigl{(}A\chi_{r_{j}}\,=\,b,A\chi_{r_{i}}\,=\,b\bigr{)},\ E]}{\Pr_{A}[\bigl{(}A\chi_{r_{i}}\,=\,b\bigr{)},\ E]}$
$=\frac{\Pr_{A}\left[\biggl{(}\bigcap\limits_{\ell\in[k]}\left(A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell}\right)\biggr{)},\ E\right]}{\Pr_{A}[\bigl{(}A\chi_{r_{i}}\,=\,b\bigr{)},\ E]}$
$=\frac{\prod\limits_{\ell\in[k]}\Pr_{A}\left[\bigl{(}A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell}\bigr{)},\ E\right]}{\Pr_{A}[\bigl{(}A\chi_{r_{i}}\,=\,b\bigr{)},\ E]}$ (3.5)
Note that $A_{\ell}$ and $b_{\ell}$ denote the $\ell$-th row of AA and $\ell$-th coordinate of $b$ respectively, and $[k]$ denotes the set $\{1,\dots,k\}$. Observe that the last equality in the above equation follows from the fact that rows of AA are independent and uniformly random samples of $\mathbb{Z}_{\Delta}^{k}$. Therefore
$\Pr_{A}\left[A_{\ell}\cdot\chi_{r_{j}}=b_{\ell}\ \mid\ A_{\ell}\cdot\chi_{r_{i}}=b_{\ell},\ E\right]=\frac{1}{\Delta^{2}}.$
Therefore, we have,
$\Pr_{A}[A\chi_{r_{j}}\,=\,b\,\mid\,A\chi_{r_{i}}\,=\,b,\ E]=\frac{\prod\limits_{\ell\in[k]}\Pr_{A}\left[A_{\ell}\cdot\chi_{r_{j}}=b_{\ell},\,A_{\ell}\cdot\chi_{r_{i}}=b_{\ell},\ E\right]}{\Pr_{A}[A\chi_{r_{i}}\,=\,b,\ E]}=\frac{1}{\Delta^{k}}.$
@@
$\Pr_{A}[\bigl{(}\exists j\neq i\text{ such that }A\chi_{r_{j}}=b\mid A\chi_{r_{i}}=b\bigr{)}\ \mid\ E]=\frac{s-1}{2s}<1.$
From the above two cases we can now say that given any $\chi_{r_{i}}\in B$, there exists a $k\times t$ matrix AA, with the entries $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$ and a $k\times 1$ matrix $b$ such that
@@
Now, let us look at the coefficients corresponding to the characters in $S$, in order to prove that the $\ell$ Fourier weight of $S\cap H$ is upper bounded by $\frac{\mu^{2}}{s}$, $H$ being the solution set of the system of linear equations $A\chi=b$, with the entries $a_{ij}\text{ are invertible in $\mathbb{Z}_{\Delta}$ for all }i\geq j$. Let us look at the expected value of $\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}$, given $A\chi_{r_{i}}=b$.
$\mathbb{E}_{A,b}\biggl{[}\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}\mid A\chi_{r_{i}}=b\ ,\ E\biggr{]}$ $=\sum_{\chi\in S}\Pr_{A,b}[\chi\in H\mid A\chi_{r_{i}}=b\ ,\ E]|\widehat{f}(\chi)|^{2}$
$\leq\frac{1}{\Delta^{k}}\times\mu^{2}\leq\frac{\mu^{2}}{2s}.$ (3.6)
@@
$\Pr_{A,b}\bigg{[}\sum_{\chi\in S\cap H}|\widehat{f}(\chi)|^{2}\geq\frac{\mu^{2}}{s}\mid A\chi_{r_{i}}=b\ ,\ E\bigg{]}\leq\frac{\frac{\mu^{2}}{2s}}{\frac{\mu^{2}}{s}}=\frac{1}{2}<1.$ (3.7)
So, if we apply union bound to Eq. 3.4 and Eq. 3.7, we get that the events $E_{1}$, $E_{2}$ and $E$ occur simultaneously with non-zero probability.
@@
Now, by Lemma 3.2, we have,
$P_{H}f(x)=\mathbb{E}_{z\in(A^{\perp})^{\perp}}[f(x-z)\chi_{r}(z)],$
which is of the form $\frac{g(\omega_{\Delta})}{\Delta^{k}}$, since by Lemma 2.27, $|(A^{\perp})^{\perp}|=\frac{|\mathbb{Z}_{\Delta}^{T}|}{|A^{\perp}|}=\Delta^{k}$ (since we have sampled the entries of the matrix AA from $\mathbb{Z}_{\Delta}$), where $g(X)\in\mathbb{Z}[X]$ is a nonzero polynomial with degree $\leq\Delta-1$. So $P_{H}f(x)$ is $k$-granular.
Also, let us look at the function $G_{1}(x)=\sum_{\chi\in S\cap H}\widehat{f}(\chi)\chi(x)$. We know, by Section 3.2, that, $\mathbb{E}_{x}G_{1}(x)^{2}\leq\frac{\mu^{2}}{s}$. So, $G_{1}(x)^{2}\leq\frac{\mu^{2}}{s}$ for some $x\in\mathbb{Z}_{p^{m}}^{n}$. For this particular $x$,